# Benchmarking Vision Language Models for Cultural Understanding

### v2

@@ 5 Benchmarking VLMs on CulturalVQA
Evaluating open-ended VQA is challenging. Traditionally, string matching has been used but it is known to underestimate model performance. Based on findings from Mañas et al. (2024), which demonstrate the effectiveness of reference-based LLM evaluation for open-ended VQA tasks, we adopt LAVE, their proposed metric, as our evaluation metric with GPT-4 as the LLM (see § F for the LLM prompt used). We validated the effectiveness of LAVE for our use case by computing correlation with human judgments. LAVE judgment agrees with human judgment 79% of the times for GPT-4, 73% of the times for Gemini, and 76% of the times for Intern-VL.
VLMs used for benchmarking
We benchmark several state-of-the-art VLMs on the proposed CulturalVQA dataset, ranging from closed-source models like GPT-4 (GPT-4o) and Gemini Pro (Gemini-Pro-Vision 1.0) to a wide variety of open-source models, ranging from 7 to 25 billion parameter count: Blip2 Li et al. (2023), InstructBlip Dai et al. (2024), Llava1.5 Liu et al. (2023), Llava_Next Liu et al. (2024), Idefics2 Laurençon et al. (2024), and Intern-VL 1.5 Chen et al. (2024). See § G for a detailed discussion of the selected models.
Open-Source Closed-Source
Country InstructBlip LLaVa1.5 Blip2 LLaVa-Next Idefics2 Intern-VL Gemini GPT-4
Brazil 10.57 40.38 32.21 45.62 54.37 52.53 66.34 76.44
Canada 17.00 50.50 58.50 62.50 69.00 67.50 65.50 72.00
China 16.52 26.09 34.78 33.04 38.26 53.04 65.22 65.22
Ethiopia 3.19 24.47 17.02 18.09 25.53 26.60 42.55 56.38
Germany 30.77 41.03 51.28 48.72 38.46 48.72 48.72 61.54
India 19.91 34.84 46.61 42.53 49.32 53.85 58.37 69.68
Iran 11.30 18.26 19.13 17.39 23.48 30.43 46.09 57.39
Nigeria 13.74 22.81 21.35 28.95 31.87 33.92 36.26 43.27
Rwanda 4.97 19.34 22.65 25.41 23.20 28.73 35.36 46.41
Turkey 21.52 24.47 33.76 33.33 37.97 41.35 56.12 59.92
USA 29.54 58.77 62.77 62.77 65.23 71.38 62.15 66.77
Average 16.27 32.81 36.37 38.03 41.51 46.18 52.97 61.36
@@
To investigate this, we employ the following baselines. LLM-only: This baseline uses an LLM to answer questions based on solely the question input. It helps gauge the extent to which the questions in our dataset can be addressed without any visual context, solely relying on the language-only cultural information encoded in the parameters of the LLM. LLM + Country: It introduces country-specific context into the LLM prompts to determine if knowing the country along with the question can already elicit the correct answer! LLM + Lens: Unlike the other two baselines, which do not rely on visual context, this baseline takes as input the image entity names extracted by Google Lens, along with the question. Thus, it helps gauge whether the questions in our dataset can be answered with only coarse-level knowledge of the visual context.
@@
We report the LAVE accuracies for zero-shot evaluation of open-source and closed-source vision-language models on the proposed CulturalVQA benchmark in Tab. 1. The average LAVE accuracy for the highest-performing model, GPT-4, is approximately 61%, with performance varying across countries from 43% to 72%. We see substantial disparity in cultural understanding across different VLMs, with the best-performing open-source model (Intern-VL for most countries) achieving an average LAVE accuracy of only 46%, and performance ranging across countries from 26% to 71%. This result indicates a considerable performance gap between closed-source models and the best-performing open-source model. This gap is particularly pronounced in countries within the African-Islamic culture (Ethiopia, Nigeria, Iran, and Turkey), with a 29.78% gap for Ethiopia, the country for which the models perform the worst.
We also conduct few-shot evaluation of VLMs to assess if solving the CulturalVQA benchmark is merely a matter of guiding the VLMs in the right direction via a handful of examples from each country. For the few-shot setup, we include one example for each country (i.e., a total of 11 examples). The few-shot prompt used in the setup is provided in § I. Our analysis in Fig. 6 reveals that few-shot prompting does not consistently improve performance over zero-shot, despite having examples from all countries. Notably, some countries like Germany, Ethiopia, and Nigeria showed clear improvements (3-8% point increase), while others such as Brazil, China, India, and Rwanda experienced performance drops or minimal gains. This inconsistency in performance improvement suggests that the effectiveness of few-shot prompting may not be uniformly beneficial across different cultural contexts. This finding implies that GPT-4’s performance on culturally nuanced tasks may be more dependent on its pre-existing knowledge base rather than the immediate context provided. These results highlight the challenging nature of our proposed CulturalVQA dataset, as it successfully captures cultural knowledge that is not easily acquired through simple few-shot prompting techniques, emphasizing the need for more advanced approaches to enhance model performance on cultural understanding tasks.
The subsequent analyses in this section are conducted on top of zero-shot results.
@@
Figure 8: Performance gap between the best open-source and closed-source models compared to human performance. Negative values indicate where models underperform relative to humans.
@@
We also calculate the Pearson correlation between human and model performance across countries. For open-source models, we observe a relatively low correlation, ranging from 0.1 to 0.4. Interestingly, for closed-source models like Gemini and GPT-4, we find a stronger correlation of 0.69 and 0.75, respectively. This suggests that the factors affecting human performance similarly influence the performance of these closed-source models. However, from Fig. 8, when comparing human and model performance using the same metric, we find that closed-source models still lag behind humans for all countries indicating that while these models follow human performance trends, there is still a marked gap in their cultural understanding compared to humans. This gap is even more pronounced for open-source models, which show an even larger discrepancy in model performance compared to human performance across all countries.
@@
How much do varying question difficulty and varying answer counts affect model performance disparity across countries?
Since we sourced questions from different annotator groups across countries, it is imperative to ask if the disparity in model performance across countries is due to the disparity in inherent difficulty levels of questions across countries. To investigate this, we analyze the correlation between the model performance and the average question length (see Fig. 3 (B) for average question length across countries). We use average question length as a proxy for question difficulty because we believe shorter questions probe more direct knowledge, whereas longer questions require more nuanced, detailed cultural knowledge. As the variance in absolute question lengths is small, looking at ranks is more meaningful, which is why we use Spearman for this analysis. We find a weak Spearman rank correlation between question length and model performance ranging between -0.3 and 0.3 for most models, with the exception of GPT-4 and Gemini, which exhibit a moderate negative correlation of -0.52 on average. Further, Fig. 11 presents a scatter plot of GPT-4 performance vs. average question lengths across countries. We can see that for most countries except Brazil, Canada, and the USA, the variance in question lengths is small. This suggests that question length is not a significant factor behind the disparity in model performance across countries.
Another factor potentially affecting the disparity in model performance is the variable number of human answers per question across countries (see Fig. 3 (C)). These human answers are used as the reference answers in the LAVE metric making it more rigid for countries with fewer references and vice-versa. To investigate this, we compute the correlation between model performance and the average number of answers per question across countries. We again use Spearman correlation since the variance in absolute values of average numbers of answers per question across countries is small. We find a very low correlation ranging between -0.3 and 0 across models, indicating that the disparity in the number of human answers does not meaningfully affect the disparity in model performance across countries.
@@
We evaluate responses from the three best-performing models, GPT-4, Gemini, and Intern-VL to questions from India, with each answer rated by 5 humans on a scale of 1 to 5, from completely correct to completely incorrect. See § K for details on the human evaluation study. Fig. 9 shows the percentage of questions that fall into each of the five scales.The models’ scores closely align with human judgments for case 1 scores, suggesting that our metric predicts answers to be correct only if they are both precise and culturally specific. We note that humans tend to rate model predictions higher than the LAVE metric. Finally, the evaluation shows that humans tend to choose the extreme ratings, considering most model responses as either fully accurate or entirely wrong.
@@
Our qualitative evaluation of the best-performing model, GPT-4, highlights its limitations in recognizing and interpreting cultural nuances. For instance, GPT-4 overlooks the cultural significance of intangible cultural concepts like coral beads in Nigeria, which symbolize wealth and heritage but are treated merely as decorative objects, as well as it fails to recognize the symbolic connection between cows and planet Earth in Indian culture (see Fig. 10). Focusing on tangible cultural concepts in Fig. 10, the model’s shortcomings are evident as it inaccurately recognizes cultural entities and objects. For instance, it mislabels Naghali, a traditional Iranian storyteller as a Dervish and mistakes a traditional Turkish tea glass for a tulip glass, commonly used for serving beer. These examples reveal how GPT-4 struggles with both tangible and intangible cultural concepts: it has difficulties distinguishing between visually similar but culturally distinct entities and objects, and it lacks a deep understanding of cultural beliefs and symbolic meanings.
@@ 6 Conclusions
In this paper, we highlight the significance of evaluating VLMs not just on general scene understanding but also on their ability to comprehend diverse cultural contexts. We introduce CulturalVQA, a novel cultural VQA benchmark for assessing VLMs on their cultural understanding. By curating a diverse collection of images from 11 countries across 5 continents and collecting 2,378 hand-crafted questions and 7,206 answers about cultural concepts presented in these images, written by annotators, we ensured a broad representation of cultural concepts pertinent to diverse cultural groups.
Benchmarking state-of-the-art models on CulturalVQA reveals notable disparities in their performance across regions. Specifically, models demonstrate substantially higher accuracy in answering questions related to North American cultures compared to African-Islamic ones. Further, we find a stark performance disparity between closed- and open-source models, with a 29.78% gap between the highest-performing proprietary model and its open-source counterpart in the country for which the models perform the worst. The benchmarked VLMs also showed varying levels of proficiency across cultural facets, performing well on questions about clothing, rituals, and traditions, but less effectively on those concerning food and drink. Our results underscore the current limitations of VLMs in achieving uniform cultural comprehension and pinpoint specific areas that require improvement.
@@ 7 Limitations
Our study faces limitations due to our data collection methods, the scope of the CulturalVQA dataset, and our focus on the English language. We approximated cultural groups using geographical regions for annotator recruitment, potentially oversimplifying cultural identities and conflating culture with nationality due to practical constraints like annotator availability. Our use of English-only data may also miss key cultural nuances available only in native languages. Although our dataset aims for cultural diversity, it does not capture the full spectrum of global cultural diversity. Future work will expand the dataset to represent diverse cultures and regions more broadly and develop multilingual datasets for greater inclusivity.
@@
Collecting culturally rich content from diverse annotators proved challenging, particularly because the images and concepts were limited to those available on English-language websites. This restriction likely omits important cultural details. Allowing annotators to skip inadequate images did not fully overcome the drawbacks of limited image quality, impacting the depth of the questions created.
@@ 8 Ethical Considerations
Despite these efforts, we recognize our approach’s limitation in equating cultural groups with national borders, potentially overlooking the complex realities of minority and diaspora communities. We urge future research to explore finer distinctions within cultural groups to enhance representation. Although we have rigorously tried to remove biases, some subjective content may persist; however, a substantial portion of the dataset has been verified as unbiased (see § E). We acknowledge these constraints but are hopeful that our work will advance the understanding of cultural nuances in VLMs.
@@ A Data Statement
All texts included in the dataset are in English, primarily authored by non-native speakers, and may thus contain ungrammatical structures both in questions and answers.
@@
All annotators come from the following 11 countries: China, Turkey, Iran, Ethiopia, Nigeria, Rwanda, Germany, USA, Canada, Brazil, and India. Other demographics such as age and gender are unknown. All annotators were compensated at an hourly rate of 10-15\$ per hour depending on a task and the number of completed HITs.

### v3

@@ 5 Evaluating VLMs on CulturalVQA
Evaluating open-ended VQA is challenging. Traditionally, string matching has been used but it is known to underestimate model performance. Based on findings from Mañas et al. (2024), which demonstrate the effectiveness of reference-based LLM evaluation for open-ended VQA tasks, we adopt LAVE, their proposed metric, as our evaluation metric with GPT-4 as the LLM (see § L for the LLM prompt used). We validated the effectiveness of LAVE for our use case by computing correlation with human judgments. LAVE judgment agrees with human judgment 79% of the times for GPT-4, 73% of the times for Gemini, and 76% of the times for Intern-VL.
VLMs used for evaluation
We benchmark several state-of-the-art VLMs on the proposed CulturalVQA dataset, ranging from closed-source models like GPT-4 (GPT-4o), Claude (Claude 3.5) and Gemini Pro (Gemini-Pro-Vision 1.0) to a wide variety of open-source models, ranging from 7 to 25 billion parameter count: Blip2 Li et al. (2023), InstructBlip Dai et al. (2024), mBLIP Geigle et al. (2023) PalliGemma Beyer et al. (2024) Llava1.5 Liu et al. (2023), Llava_Next Liu et al. (2024), Idefics2 Laurençon et al. (2024), and Intern-VL 1.5 Chen et al. (2024). See § D for a detailed discussion of the selected models.
Open-Source Closed-Source Country mBLIP LLaVa1.5 Blip2 LLaVa-Next Idefics2 Intern-VL Gemini Claude GPT-4 Brazil 25.34 40.38 32.21 45.62 54.37 52.53 66.34 66.36 76.44 Canada 38.50 50.50 58.50 62.50 69.00 67.50 65.50 66.00 72.00 China 22.61 26.09 34.78 33.04 38.26 53.04 65.22 49.57 65.22 Ethiopia 7.44 24.47 17.02 18.09 25.53 26.60 42.55 41.49 56.38 Germany 41.02 41.03 51.28 48.72 38.46 48.72 48.72 51.28 61.54 India 27.83 34.84 46.61 42.53 49.32 53.85 58.37 59.28 69.68 Iran 13.04 18.26 19.13 17.39 23.48 30.43 46.09 47.83 57.39 Nigeria 13.16 22.81 21.35 28.95 31.87 33.92 36.26 36.55 43.27 Rwanda 13.26 19.34 22.65 25.41 23.20 28.73 35.36 33.70 46.41 Turkey 28.57 24.47 33.76 33.33 37.97 41.35 56.12 51.26 59.92 USA 38.77 58.77 62.77 62.77 65.23 71.38 62.15 64.92 66.77 Average 24.50 32.81 36.37 38.03 41.51 46.18 52.97 51.66 61.36
@@
To investigate this, we employ the following baselines. LLM-only: This baseline uses an LLM to answer questions based solely on the question input. It helps gauge how well the questions can be addressed without visual context, relying only on the cultural knowledge encoded in the LLM. LLM + Country: It introduces country-specific context into the LLM prompts to determine if knowing the country along with the question can already elicit the correct answer. LLM + Lens: This baseline uses image entity names extracted by Google Lens (Google, 2017) along with the question as input, unlike the other baselines that lack visual context. It helps assess whether coarse-level visual knowledge is sufficient to answer the questions.
@@
Figure 6: Performance gap between the best open-source (one of InternVL, Idefics2, BLIP2) and closed-source models (GPT-4o) compared to human performance. Negative values indicate where models underperform relative to humans.
@@
We report the LAVE accuracies for zero-shot evaluation of VLMs on the proposed CulturalVQA benchmark in Tab. 2 and Tab. 4. The average LAVE accuracy for the highest-performing model, GPT-4, is approximately 61%, with performance varying across countries from 43% to 72%. We see substantial disparity in cultural understanding across different VLMs, with the best-performing open-source model (Intern-VL for most countries) achieving an average LAVE accuracy of only 46%, and performance ranging across countries from 26% to 71%. This result indicates a considerable performance gap between closed-source models and the best-performing open-source model. It is particularly pronounced in countries within the African-Islamic culture (Ethiopia, Nigeria, Iran, and Turkey), with a 29.78% gap for Ethiopia, the country for which the models perform the worst. We also conduct few-shot evaluation of VLMs but find that it does not significantly impact performance (see § E for more details). Hence, the subsequent analyses in this section are conducted on top of zero-shot results.
@@
Do multilingual VLMs perform better in culturally diverse settings?
One might expect that multilingual VLMs may demonstrate superior performance due to their exposure to culturally diverse data. However, our analysis of multilingual models, mBLIP and PaliGemma, on CulturalVQA reveals a more nuanced picture. From Tab. 2, mBLIP, built on top of monolingual BLIP2, consistently underperforms it despite multilingual training. This could be due to the quality of the machine-translated data used in mBLIP and the LLM backbone used (mT0 (Muennighoff et al., 2023) in mBLIP vs. FlanT5 (Chung et al., 2022) in BLIP2). Also, from Tab. 4 we observe that PaliGemma shows significant disparities across countries despite large-scale multilingual training. This is possibly due to its smaller size (3B) which suggests that multilingual data exposure alone is insufficient for cultural understanding.
@@
We also calculate the Pearson correlation between human and model performance across countries. For open-source models, we observe a relatively low correlation, ranging from 0.1 to 0.4. Interestingly, for closed-source models like Gemini and GPT-4, we find a stronger correlation of 0.69 and 0.75, respectively. This suggests that the factors affecting human performance similarly influence the performance of these closed-source models. However, from Fig. 6, when comparing human and model performance using the same metric, we find that closed-source models still lag behind humans for all countries indicating that while these models follow human performance trends, there is still a marked gap in their cultural understanding compared to humans. This gap is even more pronounced for open-source models, which show an even larger discrepancy across all countries.
@@
How much does varying question difficulty and varying answer counts affect model performance disparity across countries?
Since we sourced questions from different annotator groups across countries, it is imperative to ask if the disparity in model performance across countries is due to differences in inherent question difficulty across countries. To investigate this, we analyze the Spearman rank correlation  [^7] We use this metric as the variance in lengths is small, making rank-based analysis more meaningful. between the model performance and the average question length (see Fig. 3 (B) for average question length across countries). We use average question length as a proxy for question difficulty - assuming shorter questions probe more direct knowledge, while longer ones require nuanced cultural understanding. We found a weak correlation between question length and model performance (-0.3 to 0.3) for most models, with the exception of GPT-4 and Gemini, which showed a moderate negative correlation of -0.52 on average. As illustrated in Fig. 12, for most countries except Brazil, Canada, and the USA, the variance in question lengths is small, suggesting that question length is not a significant factor behind the disparity in model performance across countries.
Another factor potentially affecting the disparity in model performance is the variable number of human answers per question across countries (see Fig. 3 (C)). These human answers are used as the reference answers in the LAVE metric making it more rigid for countries with fewer references and vice-versa. To investigate this, we compute the Spearman correlation7 between model performance and the average number of answers per question across countries. We find a very low correlation ranging between -0.3 and 0 across models, indicating that the disparity in the number of human answers does not meaningfully affect the disparity in model performance across countries.
@@
We evaluate responses from the three best-performing models, GPT-4, Gemini, and Intern-VL to questions from India, with each answer rated by 5 humans on a scale of 1 to 5, from completely correct to completely incorrect. See § H for details on the human evaluation study. Fig. 11 shows the percentage of questions that fall into each of the five scales.The models’ scores closely align with human judgments for case 1 scores, suggesting that our metric predicts answers to be correct only if they are both precise and culturally specific. We note that humans tend to rate model predictions higher than the LAVE metric. Finally, the evaluation shows that humans tend to choose the extreme ratings, considering most model responses as either fully accurate or entirely wrong.
@@
Our qualitative evaluation of the best-performing model, GPT-4, highlights its limitations in recognizing and interpreting cultural nuances. For instance, GPT-4 overlooks the cultural significance of intangible cultural concepts like coral beads in Nigeria, which symbolize wealth and heritage but are treated merely as decorative objects, as well as it fails to recognize the symbolic connection between cows and planet Earth in Indian culture (see Fig. 8). Focusing on tangible cultural concepts in Fig. 8, the model’s shortcomings are evident as it inaccurately recognizes cultural entities and objects. For instance, it mislabels Naghali, a traditional Iranian storyteller as a Dervish and mistakes a traditional Turkish tea glass for a tulip glass, commonly used for serving beer. These examples reveal how GPT-4 has difficulties distinguishing between visually similar but culturally distinct entities and objects, and it lacks a deep understanding of cultural beliefs and symbolic meanings.
@@ 6 Conclusions
In this paper, we introduce CulturalVQA, a novel VQA benchmark for assessing VLMs on their cultural understanding. By curating a diverse collection of images from 11 countries across 5 continents and collecting 2,378 hand-crafted questions and 7,206 answers about cultural concepts presented in these images, written by annotators, we ensured a broad representation of cultural concepts pertinent to diverse cultural groups.
Benchmarking state-of-the-art models on CulturalVQA reveals notable disparities in their performance across regions. Models perform much better on North American cultures compared to African-Islamic ones. Further, we find a stark performance disparity between closed- and open-source models, with a 29.78% gap between the highest-performing closed-source and open-source models for the lowest-performing country. VLMs also show varying proficiency across cultural facets, excelling in questions about clothing, rituals, and traditions but struggling with food and drink. Our results underscore the current limitations of VLMs in achieving uniform cultural comprehension and pinpoint specific areas that require improvement.
@@ 7 Limitations
Our study faces limitations due to our data collection methods, the scope of the CulturalVQA dataset, and our focus on the English language. We approximated cultural groups using geographical regions for annotator recruitment, potentially oversimplifying cultural identities and conflating culture with nationality due to practical constraints like annotator availability. We acknowledge that some cultural concepts may lack local terms that can be effectively represented in English letters  [^8] https://tinyurl.com/3zvjsv6x. Hence, our use of English-only data may also miss key cultural nuances available only in native languages. In such cases, collecting annotations in native languages would help mitigate this issue. However, we emphasize that our benchmark, despite being in English, is already challenging enough for the models, as evidenced by the significant disparity in model performance across cultures. In § 4, our analysis revealed that 80% of the answers contain at least one word matching an English Wikipedia page, while 20% lack such a match. This suggests that these answers may be multilingual, which presents a limitation for our English-only benchmark. Although our dataset aims for cultural diversity, it does not capture the full spectrum of global cultural diversity. Future work will expand the dataset to represent diverse cultures and regions more broadly and develop multilingual datasets for greater inclusivity.
@@
Collecting culturally rich content from annotators proved challenging, particularly because the images and concepts were limited to those available on English-language websites. This restriction likely omits important cultural details. Allowing annotators to skip inadequate images did not fully overcome the drawbacks of limited image quality, impacting the depth of the questions created.
@@ 8 Ethical Considerations
Despite these efforts, we recognize our approach’s limitation in equating cultural groups with national borders, potentially overlooking the complex realities of minority and diaspora communities. We urge future research to explore finer distinctions within cultural groups to enhance representation. Although we have rigorously tried to remove biases, some subjective content may persist; however, a substantial portion of the dataset has been verified as unbiased (see § C). We acknowledge these constraints but are hopeful that our work will advance the understanding of cultural nuances in VLMs.
@@ A Data Statement
Country Brazil Canada China Ethiopia Germany India Iran Nigeria Rwanda Turkey USA No. Annotators 5 6 6 4 11 6 4 8 7 4 5
Table 3: Number of Annotators by Country
Model Brazil Canada China Ethiopia Germany India Iran Nigeria Rwanda Turkey USA Avg. PaliGemma 38.87 54.50 20.87 9.57 35.89 35.52 13.04 19.88 14.36 26.05 54.77 28.67 InstructBLIP 10.57 17.00 16.52 3.19 30.77 19.91 11.30 13.74 4.97 21.52 29.54 16.27
Table 4: Performance of InstructBLIP and PaliGemma on CulturalVQA
@@
All texts in the dataset are in English, primarily authored by non-native speakers, and may contain ungrammatical structures in both questions and answers. We build our dataset in English to disentangle multicultural understanding from multilingual comprehension.
@@
All annotators come from the following 11 countries: China, Turkey, Iran, Ethiopia, Nigeria, Rwanda, Germany, USA, Canada, Brazil, and India. Initially, we attempted to engage professional annotators from the Amazon Mechanical Turk (MTurk) platform. However, we encountered challenges in finding sufficient presence of annotators from some of the targeted countries. Therefore, we expanded our search to other communities with a broad cultural representation, including Masakhane, an African NLP organization, and Mila, an international academic AI research institute. All annotators are either natives of the country they annotated for or have resided there for at least 18 years, ensuring they have sufficient cultural context and lived experiences required for the task. We conducted multiple pilot rounds to ensure that annotators adhere to our guidelines and are fluent in English. Other demographics such as age and gender are unknown. All annotators were compensated at an hourly rate of 10-15\$ per hour depending on a task and the number of completed HITs. The number of unique annotators from each country can be found in Tab. 3.